# A Convex-Inspired Neural Construction for Structured and Generalizable Nonlinear Model Reduction

### v1

@@ Abstract.
Real-time simulation of deformable objects relies on model reduction to achieve interactive performance while maintaining physical fidelity. Traditional linear methods, such as principal component analysis (PCA), provide structured and predictable behavior thanks to their linear formulation, but are limited in expressiveness. Nonlinear model reduction, typically implemented with neural networks, offers richer representations and higher compression; however, without structural constraints, the learned mappings often fail to generalize beyond the training distribution, leading to unstable or implausible deformations. We present a symmetric, convex-inspired neural formulation that bridges the gap between linear and nonlinear model reduction. Our approach adopts an input-convex neural network (ICNN) augmented with symmetry constraints to impose structure on the nonlinear decoder. This design retains the flexibility of neural mappings while embedding physical consistency, yielding coherent and stable displacements even under unseen conditions. We evaluate our method on challenging deformation scenarios involving forces of different magnitudes, inverse directions, and sparsely sampled training data. Our approach demonstrates superior generalization while maintaining compact reduced spaces, and supports real-time interactive applications.
Reduced-order modeling, Implicit neural representation, Computational design
@@ 1. Introduction
A widely adopted paradigm for model reduction is data-driven learning. Early approaches (Barbič and James, 2005) collected snapshot data from full-space simulations and construct a linear subspace via principal component analysis (PCA). This linear approximation enables fast runtime evaluation but limits expressiveness due to the inherent linearity of the representation. To address this, Fulton et al. (Fulton et al., 2019) introduced a nonlinear model reduction framework that learns an additional mapping from a low-dimensional latent space to the PCA-reduced coordinates. This nonlinear formulation achieves greater compression, faster convergence, and improved online performance compared to its linear counterpart.
@@
While both linear and nonlinear data-driven methods perform well within the latent space spanned by their training data, their behavior beyond this range differs fundamentally. For linear bases such as PCA, the linearity of the mapping enforces a structured form of generalization: when the latent coordinates move outside the training region, the corresponding deformations scale consistently along those directions. In contrast, nonlinear model reduction—typically parameterized by neural networks—lacks such structural constraints. Outside the training distribution, the learned mapping can behave unpredictably, often producing implausible or unstable deformations. This exposes a fundamental dilemma: linear models offer structured predictability but limited expressiveness, whereas nonlinear models provide flexibility at the cost of generalization. Our work bridges this divide by introducing a formulation that preserves the structural consistency characteristic of linear models while retaining the expressive power of nonlinear ones.
In our formulation, the mapping between the reduced space and the full space remains nonlinear. The key question, then, is what structure within this nonlinear functional is desirable for model reduction of deformable objects. We introduce a symmetric, convexity-induced formulation that imposes additional structure on the displacement mapping, leading to more coherent and stable behavior compared to unconstrained nonlinear models.
To realize this formulation, we adopt the input-convex neural network (ICNN) (Amos et al., 2017), which guarantees convexity of the output with respect to the input by construction. However, convexity alone does not capture the natural symmetry of many physical systems: in elastic deformation, for instance, displacements of opposite direction often correspond to states of equivalent energy. A purely convex mapping cannot represent this behavior, as convex functions are inherently one-sided. To address this limitation, we extend the ICNN with an additional symmetry constraint that enforces antisymmetric response in the displacement field. This design embeds the intrinsic symmetry of physical systems directly into the nonlinear mapping, yielding a representation that is both structured and expressive.
@@
We evaluate our method by examining its generalization beyond the training distribution. Specifically, we test under external forces that differ substantially from those seen during training, including forces of larger magnitude and opposite direction. Across these challenging settings, our nonlinear mapping demonstrates markedly improved generalization, producing qualitatively more realistic deformations and quantitatively faster convergence.
Because our formulation remains nonlinear yet structured, it also achieves higher compression: the reduced space can be represented with fewer dimensions while maintaining accuracy. This compactness improves robustness when only a small number of cubature points are available. In addition, our symmetric convex formulation yields smoother interpolation even when the training data are sparsely sampled. Finally, our implementation achieves real-time performance, supporting interactive deformation and user manipulation.
In summary, we
  • Bridge linear and nonlinear model reduction. We address the fundamental trade-off between the structured generalization of linear models and the expressive flexibility of nonlinear ones, introducing a unified formulation that embeds linear-like structure into nonlinear mappings.
  • Propose a symmetric-induced convex neural formulation for deformable simulation. We design a structured decoder based on input-convex neural networks augmented with symmetry constraints, ensuring coherent displacement behavior and stable extrapolation beyond the training distribution.
  • Provide comprehensive evaluation of generalization and compression. We demonstrate strong generalization to unseen forces, boundary conditions, and sparsely sampled datasets, while achieving compact reduced spaces, robust cubature performance, and real-time interactive simulation.
@@ 2. Related Work
Figure 2. Our convex-inspired structure builds upon the input-convex neural network (ICNN), in which the output is, by construction, convex with respect to the input. This convexity is ensured through non-negative weights and non-decreasing activation functions.
@@ 2.1. Reduced-space Simulation using Neural Networks
Neural networks have become a popular tool for accelerating reduced-order modeling in physics-based simulation, with applications spanning fluids (Deng et al., 2023; Jain et al., 2024; Kim et al., 2019; Tao et al., 2024), contact and collisions (Cai et al., 2022; Romero et al., 2021; Yang et al., 2020), and deformable materials (Chen et al., 2023a; Feng et al., 2024; Chen et al., 2023b). While prior work has explored both linear and nonlinear formulations—such as learning linear subspaces with neural priors (Chang et al., 2023)—we specifically focus on nonlinear model reduction for deformable objects in this paper, as it forms the core of our method.
@@
Our goal is similar in that we also seek to regularize the learned latent space. Like previous methods, ours can be combined with a PCA layer. However, instead of relying on additional loss terms (Sharp et al., 2023; Lyu et al., 2024), we adopt a network architecture that is convex by design. This convexity holds even for inputs that lie outside the training sample distribution.
@@ 2.2. Regularization of Neural Networks
In reduced-space simulation, decoder overfitting can lead to nonphysical artifacts, slow convergence during energy minimization, and poor generalization to out-of-distribution inputs. Previous approaches have attempted to mitigate these issues through loss-based regularization (Sharp et al., 2023; Lyu et al., 2024), but these methods rely on explicit penalty terms. This strategy has two key limitations. First, regularization is only effective if the loss is sufficiently optimized, making it difficult to determine the optimal training epoch. Second, these approaches require sampling the latent space; however, because the latent space is typically unbounded, it is impossible to samplenex exhaustively. As a result, the desired regularization properties may not hold for latent codes outside the sampled region. In contrast, our formulation enforces convexity everywhere in the latent space, providing stronger and more reliable guarantees.
We enforce regularization architecturally via Input Convex Neural Networks (ICNNs) (Amos et al., 2017). The convex‐by‐design decoder admits a globally smooth latent energy landscape without any extra penalty terms, leading to more stable optimization, faster convergence, and stronger generalization in reduced‐space simulation.
Figure 3. We demonstrate the advantage of input convex neural networks (ICNNs) on the convex function $z=x^{2}+y^{2}$. Both an ICNN and a standard MLP are trained on samples within a disk of radius 2 (dashed blue circle). The left plot shows the ground truth. The middle plot shows the ICNN prediction, which accurately captures the convexity and generalizes smoothly outside the training region. The right plot shows the MLP prediction, which fits well inside the training region but produces non-convex distortions outside.
3. Background: Input Convex Neural Network and Nonlinear Model Reduction
3.1. Input Convex Neural Network
An Input Convex Neural Network (ICNN) (Amos et al., 2017) is a feed-forward architecture $f_{\theta}:\mathbb{R}^{k}\to\mathbb{R}^{N}$ whose output is convex with respect to its input $\mathbf{q}\in\mathbb{R}^{k}$. The network is constructed as a sequence of activations $\mathbf{z}_{0},\mathbf{z}_{1},\dots,\mathbf{z}_{m}$, where each layer is defined by
(1) $\mathbf{z}_{i+1}=g_{i}\left(\mathbf{W}_{i}^{(z)}\mathbf{z}_{i}+\mathbf{W}_{i}^{(q)}\mathbf{q}+\mathbf{b}_{i}\right),\quad\text{for }i=0,\dots,m-1,$
with output $f_{\theta}(\mathbf{q})=\mathbf{z}_{m}$. Here, $g_{i}$ denotes the element-wise activation function, and the trainable parameters are
$\theta=\{\mathbf{W}_{i}^{(q)},\mathbf{W}_{i}^{(z)},\mathbf{b}_{i}\}_{i=0}^{m-1}.$
The initial activation is set as $\mathbf{z}_{0}=0$, and the first skip weight is fixed to zero, i.e., $\mathbf{W}_{0}^{(z)}\equiv 0$, to ensure convexity from the first layer onward.
Convexity of the overall function $f_{\theta}(\mathbf{q})$ is guaranteed under two conditions: (1) all weights $\mathbf{W}_{i}^{(z)}$ for $i\geq 1$ are constrained to be element-wise non-negative, and (2) all activation functions $g_{i}$ are convex and non-decreasing (e.g., ReLU or softplus). These architectural constraints ensure that each layer preserves convexity, resulting in a globally convex function with respect to the input. We refer readers to (Amos et al., 2017) for a detailed proof of the convexity of this formulation.
@@ 3.2. Nonlinear Model Reduction
Let $\mathbf{x}_{0}\in\mathbb{R}^{N}$ denote the rest configuration of a discretized mesh with $n$ nodes and $m$ elements, where $N=3n$ in the three-dimensional case. At time $t$, the deformation is described by a displacement vector $\mathbf{u}\in\mathbb{R}^{N}$, such that the current vertex positions are given by
@@ 3.3. Learning the Configuration Mapping
To construct the nonlinear mapping $f$, previous works (Fulton et al., 2019; Sharp et al., 2023) propose to parameterize $f$ with a multilayer perceptron (MLP) possessing trainable weights $\theta$. Let $f_{\theta}:\mathbb{R}^{k}\to\mathbb{R}^{N}$ denote this mapping, and $g_{\theta}:\mathbb{R}^{N}\to\mathbb{R}^{k}$ its encoder counterpart. Given a collection of observed system configurations $Q=\{\mathbf{u}_{i}\in\mathbb{R}^{N}\}$, the autoencoder framework is employed to learn a compact latent representation that captures the intrinsic structure of the configuration manifold.
The parameters $\theta$ are obtained by minimizing the reconstruction loss
(5) $\min_{\bm{\theta}}\frac{1}{|\mathcal{B}|}\sum_{\mathbf{u}_{i}\in\mathcal{B}}\|f_{\theta}(g_{\theta}(\mathbf{u}_{i}))-\mathbf{u}_{i}\|_{M}^{2},$
where $\mathcal{B}\subset Q$ is a randomly drawn mini-batch and $|\mathcal{B}|$ its cardinality. Here, $\|\cdot\|_{M}$ denotes the norm induced by the manifold metric, ensuring that reconstruction fidelity is measured consistently with the underlying geometric structure of the data.
Minimizing Eq. (5) yields a decoder $f_{\theta}$ that provides a smooth, differentiable mapping from the latent coordinates to the configuration space, thereby enabling nonlinear model reduction on manifolds of complex deformations.
4. Regularizing Neural Subspace with ICNN

### v2

# Odd-DC: Generalizable Neural Model Reduction via Odd Difference-of-Convex Structure

@@ Abstract.
Model reduction is essential for real-time simulation of deformable objects. Linear techniques such as PCA provide structured and predictable behavior, but their limited expressiveness restricts accuracy under large or nonlinear deformations. Nonlinear model reduction with neural networks offers richer representations and higher compression; however, without structural constraints, the learned mapping from latent coordinates to displacements often generalizes poorly beyond the training distribution.
We present an odd difference-of-convex (DC) neural formulation that bridges linear and nonlinear model reduction. Our goal is to obtain a latent space that behaves reliably under unseen load magnitudes and directions. To improve extrapolation in magnitude, we introduce convexity into the decoder to discourage oscillatory responses. Yet convexity alone cannot represent the odd symmetry required by many symmetric systems, which is crucial for generalization to inverse force directions. We therefore adopt a DC formulation that preserves the stabilizing effect of convexity while explicitly enforcing odd symmetry. Practically, we realize this structure using an input-convex neural network (ICNN) augmented with symmetry constraints.
Across challenging deformation scenarios with varying magnitudes and reversed load directions, our method demonstrates stronger generalization than unconstrained nonlinear reductions while maintaining compact latent spaces and real-time performance. Our DC formulation extends to both mesh-based and neural-field reductions, demonstrating applicability across multiple classes of neural nonlinear model reduction.
Reduced-order modeling, Implicit neural representation, Dimensionality reduction and manifold learning
@@ 1. Introduction
A widely adopted paradigm for model reduction is data-driven method. Early approaches (Barbič and James, 2005) collected snapshot data from full-space simulations and constructed a linear subspace via principal component analysis (PCA). This linear approximation enables fast runtime evaluation but limits expressiveness due to the inherent linearity of the representation. To address this, Fulton et al. (Fulton et al., 2019) introduced a nonlinear model reduction framework that learns an additional mapping from a low-dimensional latent space to the PCA-reduced coordinates. This nonlinear formulation achieves greater compression, faster convergence, and improved online performance compared to its linear counterpart.
@@
While both linear and nonlinear data-driven methods perform well within the latent space spanned by their training data, their behavior beyond this range differs fundamentally. For linear bases such as PCA, the linearity of the mapping enforces a structured form of generalization: when the latent coordinates move outside the training region, the corresponding deformations scale consistently along those directions. In contrast, nonlinear model reduction—typically parameterized by neural networks—lacks such structural constraints. Outside the training distribution, the displacements can behave unpredictably, often producing implausible or unstable deformations. This exposes a fundamental dilemma: linear models offer structured predictability but limited expressiveness, whereas nonlinear models provide flexibility at the cost of generalization. Our work bridges this divide by introducing a formulation that preserves the structural consistency characteristic of linear models while retaining the expressive power of nonlinear ones.
In nonlinear model reduction, the mapping from reduced coordinates to full-space displacements remains highly expressive. The key question is therefore what structure should this mapping possess to generalize beyond the training set. Our goal is robustness to unseen load magnitudes and directions. To promote reliable extrapolation in magnitude, we seek a latent-to-displacement map whose response grows coherently rather than oscillating, motivating the introduction of convexity as a structural prior. Unlike loss-based regularizers that act only on sampled data, convexity constrains the mapping globally, including unsampled regions of the latent space that are inevitably visited during extrapolation.
However, convexity alone is insufficient for many approximately symmetric mechanical systems. In a wide range of elastic settings, loads applied in opposite directions tend to induce approximately opposite displacements. A purely convex mapping cannot automatically express such behavior, which can limit extrapolation when forces change sign. We therefore introduce an odd-symmetric difference-of-convex (DC) formulation, inspired by classical DC theory (Hartman, 1959), which enables antisymmetric responses while retaining the stabilizing role of convexity. This structure targets generalization in both magnitude (through convexity) and direction (through approximate odd symmetry).
To realize this DC construction in practice, we implement each convex component with an input-convex neural network (ICNN) (Amos et al., 2017), ensuring convexity with respect to the latent coordinates by design. We then induce odd symmetry by evaluating the network at a latent code and at its sign-reversed counterpart and combining the results in an antisymmetric manner. This strategy preserves the convex structure of the trainable network while producing a decoder that responds with opposite displacements to opposite latent directions.
@@
We evaluate our method by focusing on generalization beyond the training distribution. Specifically, we test under external forces that differ substantially from those seen during training, including larger magnitudes and reversed load directions across more than fifty deformation scenarios. In these challenging settings, our approach demonstrates stronger generalization than unconstrained nonlinear reductions, producing qualitatively more realistic deformations and quantitatively improved accuracy while maintaining compact latent spaces and real-time performance.
Because our formulation remains nonlinear yet structured, it enables higher compression: accurate behavior can be achieved with fewer reduced dimensions, improving robustness when only a small number of cubature points are available. The proposed DC structure is also representation-agnostic. We show that it can be integrated into both mesh-based nonlinear reductions and continuous neural-field–based reductions, demonstrating applicability across multiple classes of neural model reduction while supporting real-time, interactive deformation.
In summary, we proposed
  • A Difference-of-Convex structural prior for nonlinear model reduction. We propose an odd-symmetric DC decoder that improves extrapolation to unseen magnitudes and inverse directions while enabling higher compression and robust cubature, validated on 50+ mesh-based scenarios.
  • Extension to continuous model reductions. We show that the same DC prior regularizes continuous ROMs, outperforming prior discretization-agnostic nonlinear methods and enabling training and reconstruction from real-world data with generalization.
Figure 3. Effect of our odd-symmetric construction. Our odd-function formulation models the intrinsic symmetry of the system, enabling better generalization to deformations with opposite displacements. We train the model on a single deformation trajectory and visualize the resulting displacements from both the training latent codes and their inverses. Without the odd-function term (left), the model produces plausible deformations for training codes but exhibits higher variance and artifacts under latent inversion. With the odd-symmetric formulation (right), the deformations remain consistent across both original and inverted latent codes, demonstrating improved symmetry and generalization.
@@ 2.1. Reduced-space Simulation using Neural Networks
Neural networks have become a popular tool for accelerating reduced-order modeling in physics-based simulation, with applications spanning fluids (Deng et al., 2023; Jain et al., 2024; Kim et al., 2019; Tao et al., 2024), contact and collisions (Cai et al., 2022; Romero et al., 2021; Yang et al., 2020), and deformable materials (Chen et al., 2023a; Feng et al., 2024; Chen et al., 2023b). While prior work has explored both linear and nonlinear formulations—such as learning linear subspaces with neural priors (Chang et al., 2023; Modi et al., 2024)—we specifically focus on nonlinear model reduction for deformable objects in this paper, as it forms the core of our method.
@@
Our goal is similar in that we also seek to regularize the learned latent space. Like previous methods, ours can be combined with a PCA layer. However, instead of relying on additional loss terms (Sharp et al., 2023; Lyu et al., 2024), we adopt a network architecture built from convex neural components arranged in a structured Difference-of-Convex formulation, embedding regularization directly into the model design.
@@ 2.2. Regularization of Neural Networks
In reduced-space simulation, decoder overfitting can lead to nonphysical artifacts, slow convergence during energy minimization, and poor generalization to out-of-distribution inputs. Previous approaches have attempted to mitigate these issues through loss-based regularization (Sharp et al., 2023; Lyu et al., 2024), but these methods rely on explicit penalty terms. This strategy has two key limitations. First, regularization is only effective if the loss is sufficiently optimized, making it difficult to determine the optimal training epoch. Second, these approaches require sampling the latent space; however, because the latent space is typically unbounded, it is impossible to sample exhaustively. As a result, the desired regularization properties may not hold for latent codes outside the sampled region. In contrast, our formulation enforces convexity in the neural network, providing stronger and more reliable guarantees.
We enforce regularization architecturally via a specially designed Difference-of-Convex function, inspired by classical DC theory (Hartman, 1959). By constructing the decoder as a structured combination of convex components, we constrain its curvature and suppress highly oscillatory behavior without relying on extra penalty terms. This architectural constraint provides global regularization guarantees that hold across the entire latent space. Moreover, when combined with symmetric construction, the DC formulation naturally supports antisymmetric mappings, further enlarging the regime of reliable generalization.
3. Background: Nonlinear Model Reduction and Generalization
@@ 3.1. Nonlinear Model Reduction
Let $\mathbf{x}_{0}\in\mathbb{R}^{N}$ denote the rest configuration of a discretized mesh with $n$ nodes and $m$ elements, where $N=dn$ with $d\in\{2,3\}$ denoting the spatial dimension of the system. At time $t$, the deformation is described by a displacement vector $\mathbf{u}\in\mathbb{R}^{N}$, such that the current vertex positions are given by
@@ 3.2. Learning the Configuration Mapping
To construct the nonlinear mapping $f$, previous works (Fulton et al., 2019; Sharp et al., 2023) propose to parameterize $f$ with a multilayer perceptron (MLP) possessing trainable weights $\theta$. Let $f_{\theta}:\mathbb{R}^{k}\to\mathbb{R}^{N}$ denote this mapping, and $E_{\theta}:\mathbb{R}^{N}\to\mathbb{R}^{k}$ its encoder counterpart. Given a collection of observed system configurations $Q=\{\mathbf{u}_{i}\in\mathbb{R}^{N}\}$, the autoencoder framework is employed to learn a compact latent representation that captures the intrinsic structure of the configuration manifold.
The parameters $\theta$ are learned by minimizing the reconstruction loss
(4) $\min_{\bm{\theta}}\frac{1}{|\mathcal{B}|}\sum_{\mathbf{u}_{i}\in\mathcal{B}}\|f_{\theta}(E_{\theta}(\mathbf{u}_{i}))-\mathbf{u}_{i}\|_{M}^{2},$
where $\mathcal{B}\subset Q$ is a mini-batch of training configurations. The encoder $E_{\theta}$ is used only during training to infer latent coordinates that best reconstruct each state $\mathbf{u}_{i}$, while the decoder $f_{\theta}$ provides the mapping employed at test time. The norm $\|\cdot\|_{M}$ is induced by the manifold metric, ensuring that reconstruction accuracy is measured consistently with the geometric structure of the data.
Minimizing Eq. (4) thus yields a decoder that maps reduced coordinates to the full space, enabling nonlinear model reduction for complex deformation manifolds.
4. Regularizing Neural Subspace with Difference-of-Convex Function